Find the maximum and minimum values of the function f(x,y)=2x^2+3y^2−4x−5 on the domain x2+y2≤225.

What is the maximum value of f(x,y)?


List the point(s) where the function attains its maximum as an ordered pair, such as (-6,3), or a list of ordered pairs if there is more than one point, such as (1,3), (-4,7).

What is the minimum value of f(x,y) ?


List points where the function attains its minimum as an ordered pair, such as (-6,3), or a list of ordered pairs if there is more than one point, such as (1,3), (-4,7).
1
Expert's answer
2020-04-08T16:28:28-0400

Given "f(x,y) = 2x^{2}+3y^{2}-4x-5".

We need to use Lagrange multipliers for the boundary and partial derivatives to determine critical points from the interior. Consider the interior first. Then we have to solve the equations,

"\dfrac{\partial f}{\partial x} = 0\\" and "\dfrac{\partial f}{\partial y} = 0". From which we get,

"4x-4=0" and "6y=0." Thus, "(x, y) = (1,0)."

This critical point lies in the region "x^2+y^2<225" and the value of f at this point is "f(1,0)=2-4-5 = -7."


On the boundary, we find the extreme values of "f(x,y) = 2x^{2}+3y^{2}-4x-5" subject to the constraint "g(x,y) = x^{2}+y^{2}-225" using Lagrange's multipliers.


Let

"F = f + \lambda g \\= 2x^2+3y^2-4x-5+\lambda(x^2+y^2-225)"


To find the critical points we have to solve the equations,

"\dfrac{\partial F}{\partial x} = 0, 
\dfrac{\partial F}{\partial y} = 0, 
\dfrac{\partial F}{\partial \lambda} = 0\\".

"4x-4+2x\lambda = 0\\
6y+2y\lambda = 0\\
x^2+y^2 = 225"


From the second equation we get, "y=0" or "\lambda = -3". If "y = 0" , then "x = \pm15" , from the constraint.

When "\lambda = -3" , we get "x = -2" and from constraint we get, "y = \pm\sqrt{221}".

The critical points are "(1,0), (-15,0),(15,0), (-2,-\sqrt{221}), (-2,\sqrt{221})".

The values at these points are,

"f(1,0) = -7\\
f(-15,0) = 505\\
f(15,0) = 385\\
f(-2,-\sqrt{221}) = 674\\
f(-2,\sqrt{221}) = 674\\"


Therefore, the maximum value is 674 and the minimum value is -7.


The maximum value is attained at the points "(-2, -\sqrt{221}), (-2,\sqrt{221})" and the minimum value is attained at the point "(1,0)."